A girl stands 400 ft. from a tree. With her eye at ground level she sights along a yard stick and finds the top of the yard stick coincides with the top of the tree when the bottom of the yard stick is 4 feet from her eye. Find the height of the tree.
Answer by checkley71(8403)

3/4=x/400
4x=1200
x=1200/4
x=300 feet to the top of the tree.
proof
draw a triangle with the base=400, the distance from the eye to the bottom of the yard stick=4 feet & the height of the yard stick=3 feet.
now you have a 3 by 4 right triangle and an x by 400 triangle.
these are similiar triangles whose sides are proportional:
3:4, 300:400